# Inference in Tightly Identified and Large-Scale Sign-Restricted SVARs

### 24 Apr 2026 v1

@@ 4.2 Imposing Identifying Restrictions
In this section, we show how to impose identifying restrictions on the parameters using continuously differentiable transformations, which result in an unrestricted parameter vector. Our sampler runs on this vector, which means that rejection sampling is not needed to impose the identifying restrictions. Implementation of exclusion restrictions is straightforward in that no Jacobian adjustment is needed, whereas restrictions based on inequalities do require it as they are based on a change of variables. As a different set of identifying restrictions is considered in each application, there are no general transformation and Jacobian that would cover all possible cases, but we discuss a few building blocks which can be used to construct frequently encountered restrictions.
@@
This is the density that our algorithm samples from. The exact cost of computing $|J|$ will vary depending on the number and type of restrictions imposed on the model, but in applications involving sign restrictions or elasticity bounds it is negligible in comparison to the cost of evaluating the normal-inverse Wishart density. For example, if we restrict the sign of every parameter in $B$, evaluation of $\log|J|$ collapses to computing the sum over the $N^{2}$ elements of $\theta$ corresponding to $B$.
Let us consider the implementation of frequently encountered types of inequality restrictions. Although the presentation is not exhaustive, it should give an idea of how to proceed in an empirical application. Suppose first that we have a density $p_{\alpha}(\alpha)$ over a parameter $\alpha\in\mathbb{R}_{+}$, which could, for example, be an impact response parameter from the matrix $B$. A Metropolis-type sampler running on $\alpha$ produces proposals $\alpha<0$, which are always rejected, and this decreases efficiency. Instead, we can reparameterize the sampler in terms of an unrestricted parameter, so that we do not encounter a priori known regions of zero probability. For the sign restricted case, we introduce a parameter $\tilde{\alpha}\in\mathbb{R}$, and set $\alpha=e^{\tilde{\alpha}}$. The Jacobian of this transformation is $e^{\tilde{\alpha}}$, and the corresponding density function is then $p(\tilde{\alpha})=p_{\alpha}(e^{\tilde{\alpha}})e^{\tilde{\alpha}}$ by a change of variables. If we want to restrict $\alpha<0$, we set $\alpha=-e^{\tilde{\alpha}}$, and the corresponding density is $p(\tilde{\alpha})=p_{\alpha}(-e^{\tilde{\alpha}})e^{\tilde{\alpha}}$.
Another useful transformation is the logit transformation for a lower and upper bounded scalar. Suppose we want to sample a parameter $\alpha\in(a,b)$, say, an elasticity parameter which needs to be positive and also constrained from above. To that end, we define an $\tilde{\alpha}\in\mathbb{R}$, and obtain $\alpha$ as
@@
To illustrate these methods in practice, let us consider a two-variable model with a single lag and no constant. Suppose that the variables are quantity and price, and that the supply and demand shocks are identified using sign restrictions on the impact responses of the variables. The responses have opposite signs for the supply shock and the same sign for the demand shock. If the first shock is the supply shock, the first column of the impact matrix $B$ contains one positive and one negative element, whereas the second column contains two positive elements.
@@ 4.3 Estimation Algorithm
As the density function (9) does not facilitate direct sampling, we employ an MCMC method. Specifically, we use the No-U-Turn Sampler (NUTS) of [hoffman2014], which is a variant of the Hamiltonian Monte Carlo (HMC) algorithm. An HMC algorithm explores the target distribution by simulating Hamiltonian dynamics, which allows the sampler to construct long trajectories while yielding a high acceptance probability by utilizing the gradients of the target log-density. At a high level, the HMC sampler can be thought of as a Metropolis-Hastings algorithm with a complicated proposal distribution. Its performance depends on tunable parameters $\epsilon,L$ and $M$. Starting from the current draw, the HMC sampler simulates a trajectory by integrating Hamiltonian dynamics using $L$ steps of size $\epsilon$ with a mass matrix $M$, whose optimal choice is the inverse of the covariance matrix of the target posterior distribution [betancourt2018, Section 4.3]. The resulting endpoint is then used as a proposal.
The NUTS solves the issue of tuning $L$ by dynamically expanding the path taken within each iteration until it starts to make a U-turn towards the initial state. The implementation of NUTS in Stan further provides routines for adapting the step size and metric matrix during warmup, and calculates the gradient of the target log-density by automatic differentiation. These properties reduce the scope of implementation error and ensure that sampling can be done efficiently in a wide variety of situations. Further details on the HMC sampler and NUTS can be found in Appendix C.
@@
To demonstrate that an HMC sampler can effectively explore the posterior under set identification, we revisit a simulation experiment that [kitagawa2025] considered to illustrate the opposite. Although our primary interest lies in VAR models, we adopt this example because it features a known and simple posterior geometry, which allows for a transparent diagnosis of convergence issues. Specifically, we generate data from a normal distribution with mean $\mu=\sum_{i=1}^{k}\mu_{i}$ and variance unity, where the parameters $\mu_{i}$, $i=1,\dots,k$, are not identified for $k\geq 2$. We study the performance of Hamiltonian Monte Carlo in this setting for $k=2$ and k=1,000k=1{,}000, with the latter case reflecting the high-dimensional parameter spaces encountered in VAR applications. For $k=2$ and k=1,000k=1{,}000, we generate 1,000 observations from $N(2,1)$ and $N(10,1)$ distributions, respectively. The larger mean in the latter case reflects the greater number of additive components.
The left panel of Figure 1 presents a trace plot based on 12,000 draws for the case $k=2$, of which 2,000 are warmup draws. Following [kitagawa2025], the warmup draws are included. The right panel depicts the histograms of the obtained sample (without the warmup iterations) together with the analytical marginal densities, when the support of each $\mu_{i}$ is bounded between 10-10 and 10.444In the case $k=2$, the marginal density is proportional to $\Phi(\sqrt{n}(\mu_{i}+10-\bar{x}))-\Phi(\sqrt{n}(\mu_{i}-10-\bar{x}))$, where $\Phi$ is the standard normal CDF, $n$ is the sample size, and $\bar{x}$ is the sample mean. For n=1,000n=1{,}000 this density is approximately uniform. The plots indicate that the NUTS (as implemented in Stan) is able to traverse the entire posterior distribution of $\mu$ without apparent difficulties and produces draws that closely match the analytical marginal densities. This contrasts with the behavior documented by [kitagawa2025] for a standard HMC algorithm in the same set-identified setting, where the sampler failed to adequately explore the parameter space associated with $(\mu_{1},\mu_{2})$.
@@
The NUTS performs well even in a high-dimensional setting, as illustrated in Figure 2 that presents a scatter plot of the first two components of $\mu$ and the estimated autocorrelation functions for these two parameters when k=1,000k=1{,}000. The draws are dispersed without visible clustering, and the autocorrelations decay rapidly toward zero. Corresponding plots for other pairs of parameters look essentially identical. These results indicate that, even in this high-dimensional setting, the NUTS produces high-quality samples and efficiently explores the full target distribution.
@@
As [bacchiocchi2025] point out, besides flat posterior regions, a related but distinct challenge can arise when the identified set is multimodal because separated regions of posterior mass are difficult for MCMC algorithms (including HMC) to explore effectively. However as noted by [kitagawa2025], such local identification arises only in the case of non-homogenous restrictions, that is, when a parameter is set to a non-zero value. As we preclude such restrictions, multimodality does not arise, and hence the sampler does not need to traverse between distinct modes. This absence of multimodality follows from the fact that the likelihood depends on $B$ only through $BB^{\prime}$ and is therefore invariant under orthogonal rotations, implying a continuum of observationally equivalent factorizations rather than isolated peaks. [^5] To eliminate any remaining reflection symmetry, we impose the normalization $\det(B)>0$ and restrict attention to a single normalized representation of the structural parameters. This normalization does not restrict the identified set beyond our existing identifying assumptions. To see this, consider the case where at most N1N-1 columns of $B$ are identified, so the last column is left unrestricted. Let $B_{1}$ satisfy the identifying restrictions and have $\det(B_{1})<0$. Define $R=\mathrm{diag}(1,\ldots,1,-1)$ and $B_{2}=B_{1}R$, i.e., $B_{2}$ equals $B_{1}$ with the sign of its last column flipped. Then $\det(B_{2})=\det(B_{1})\det(R)=-\det(B_{1})>0$, while the first N1N-1 columns of $B_{2}$ and $B_{1}$ are equal. Hence, $B_{2}$ satisfies the same identifying restrictions and has the same economic interpretation. Moreover, the map $B\mapsto BR$ is bijective with inverse $B\mapsto BR^{-1}=BR$, which guarantees that there always exists exactly one such $B_{2}$. As the preceding example indicates, the HMC sampler is well suited for exploring such distributions.
@@
The $\hat{R}$ measures chain mixing by comparing within- and between-chain variances. For a well-mixed chain, it approaches unity as the number of draws increases. When running a single chain (as is typical in the SVAR literature), $\hat{R}$ is calculated by splitting the chain into two halves. The idea of the bulk and tail ESS is to measure sampling efficiency by adjusting the number of iterations for the autocorrelation in the sample. The bulk ESS estimates the number of effectively independent samples in the MCMC output for estimating the mean or the median of the posterior, while the tail ESS measures sampling efficiency in the tails of the distribution. [vehtari2021] recommend accepting an MCMC sample only if $\hat{R}<1.01$ and the ESS is at least 100 for a single chain. These thresholds ensure sufficient mixing and a sufficiently low Monte Carlo standard error for estimation of the quantities of interest.
@@ 5 Empirical Applications
This section revisits two empirical applications to showcase the benefits of our approach: In Section 5.1, we identify three shocks in the small-scale global oil market model of [kilian2014], and in Section 5.2, we consider the large-scale U.S. economy model recently analyzed by [chan2025]. In each case, the empirical findings coincide with prior results based on the same identifying restrictions, but we highlight the advantages of our method over existing alternatives. Also [arias2025elliptical] consider these applications to illustrate the performance of their Gibbs sampler, which facilitates straightforward comparison. As their approach represents the current state of the art for inference under identifying sign restrictions, it provides a natural benchmark for our analysis. In both applications, our method achieves convergence approximately two to eight times faster than theirs, as measured by the time required for the $\hat{R}$ statistic to fall below the conventional threshold. The obtained samples also exhibit lower autocorrelations, as reflected in greater bulk and tail ESS values in both cases. In the large-scale model, our approach is also significantly faster than the algorithm of [chan2025], although they use the asymmetric conjugate prior [[, see]]chan2022 rather than the normal-inverse Wishart prior. In Section 5.2, we also show that our approach accommodates exclusion restrictions at no additional computational cost.
@@ 5.1 Small Oil Market Model
To illustrate how our approach handles contemporaneous and dynamic sign restrictions, as well as elasticity restrictions, we estimate the oil market model of [kilian2014] on monthly data from 1973:M2 to 2009:M8.777We obtain the data from the replication package for [kilian2014] in the Journal of Applied Econometrics Data Archive. It contains four variables, namely, the change in global oil production, a measure of global real activity, the real price of oil, and the changes in global oil inventories, as well as an intercept, and monthly dummy variables. Following [kilian2014], the lag length is set at 24. The prior is flat and proportional to $|\Sigma|^{-(N+1)/2}$, which results in a conjugate normal-inverse Wishart posterior for the reduced-form parameters. Following the identification strategy of [kilian2014], we identify the flow supply, flow demand, and speculative demand shocks. In particular, we impose sign and elasticity restrictions on the impact responses, as well as dynamic sign restrictions, as summarized in Table 1. These restrictions result in a tightly identified model, and thus random sampling of the orthogonal rotation matrix $Q$ would result in extremely few accepted posterior draws, as discussed in Section 3.
@@
where $\tilde{D}$ is a $4\times 11$ matrix of dummy coefficients, with each column corresponding to a dummy variable. These parameters are transformed into an unconstrained vector $\theta$ using the transformations described in section 4.2, specifically the exponential and inverse logit transformations on $B$ and $\Psi_{1},\dots,\Psi_{12}$. We run 2,000 warmup iterations, followed by 100,000 sampling iterations.
@@
We now turn to the properties of our sampler. Figure 4 presents trace plots of the unnormalized posterior log-density and a parameter ($B_{11}$) from the impact matrix $B$. For clarity, only the first 4,000 iterations, including warmup, are shown. As can be seen, both the log-density and the parameter oscillate around a constant value starting from the first iteration onward. This indicates that our method for selecting initial values by direct sampling works well and initializes the chain near the typical set. The traceplots do not exhibit problematic behavior such as nonstationarity or high autocorrelation after warmup. Furthermore, the variance of $B_{11}$ is lower during the early warmup phase than during sampling, which reflects the dynamic adaptation of the step size and inverse metric. This adaptation leads to more efficient sampling after warmup. The remaining structural parameters exhibit similar behavior and their traceplots are therefore not shown.
@@
To compare our approach with the elliptical slice sampler of [arias2025elliptical], we implemented their sampler in R and, following their lead, save every tenth draw from the resulting chain. We run this sampler for a total of four million iterations. [^8] We also implemented the sampler in Matlab, and it turned out faster than the R implementation. All results involving runtimes are based on this faster implementation. To assess convergence, we use the $\hat{R}$ statistic (see Section 4.3), calculated from a single chain by considering samples of increasing length, with warmup iterations included. [^9] Because there is no clear way to ensure that the discarded warmup samples are comparable across samplers, we include the warmup iterations in the analysis. As recommended by [vehtari2021], we interpret values of $\hat{R}$ above 1.01 as an indication of potential convergence issues.
@@
Figure 5 plots the maximum values of the $\hat{R}$ statistic for samples obtained using both samplers. The reported sample size is increased in increments of 1,000 iterations. For the elliptical slice sampler, the maximum $\hat{R}$ is the maximum over the elements of $\Sigma$ and $Q$ , whereas for the NUTS it is the maximum over the elements of BB. [^10] In our experience, the unidentified parameters $B$ and $Q$ mix most slowly. To reduce computational cost, we therefore do not compute $\hat{R}$ for all parameters. As can be seen in the figure, $\hat{R}$ converges in fewer iterations when using the NUTS. For the elliptical slice sampler, $\hat{R}$ first falls below the threshold after approximately 29 thousand iterations (corresponding to roughly 14 minutes), but requires about 100 thousand saved draws (47 minutes) to remain consistently below the threshold of 1.01. Furthermore, it exceeds the threshold again at around 180,000 iterations (84 minutes). In contrast, the NUTS reaches the 1.01 threshold after roughly eight thousand iterations, corresponding to a runtime of 28 minutes.
@@
Figure 5: The maximum $\hat{R}$ statistic for samples of different sizes from the elliptical slice sampler and NUTS. Horizontal dashed lines are located at 1.00 and 1.01. $\hat{R}$ values are calculated for the elements of $B$ for the NUTS, and the elements of $\Sigma$ and $Q$ for the elliptical slice sampler. First four thousand iterations not plotted for legibility.
@@
To investigate the sampling efficiency of the samplers, we consider the bulk and tail ESS measures of [vehtari2021]. In line with the $\hat{R}$, they are calculated for all elements of $B$ (for the NUTS), and $\Sigma$ and $Q$ (for the slice sampler), again using chains of increasing length and including warmup iterations. Figure 6 presents the minimum ESS values over these parameters for the NUTS and the elliptical slice sampler. As can be seen, the NUTS is more efficient in terms of ESS per iteration under both measures. This is especially evident in the case of the bulk ESS, indicating that our approach yields more reliable estimates of posterior means for a given number of iterations. Importantly, for both methods, ESS increases approximately linearly with the number of iterations and reaches levels at which the Monte Carlo error is negligible for inference.
@@
Figure 6: The minimum of bulk and tail ESS statistics for samples of different sizes of the elliptical slice sampler and NUTS. The ESS values are calculated for elements of $B$ for the NUTS, and $\Sigma$ and $Q$ for the elliptical slice sampler.
@@
The non-monotone behavior of the $\hat{R}$ statistic and the lower ESS values for the elliptical slice sampler are likely due to the high autocorrelation in the sampled chain. Figure 7 presents the autocorrelation function for $B_{11}$ from the NUTS (top panel), and $Q_{11}$ from the elliptical slice sampler (bottom panel). The latter exhibits substantially higher autocorrelation, which reduces sampling efficiency. Similar patterns are observed for the other elements of the matrices $B$ and $Q$ in this application.
We also assess convergence using the minimum multivariate ESS, as suggested by [vats2019]. To achieve a precision of 0.05 for the 95% confidence region our approach requires 63 minutes of sampling, whereas the elliptical slice sampler requires 29 minutes. Precision of 0.05 implies that the Monte Carlo error is 5% of the variability in target distribution. While this metric suggests that the elliptical slice sampler is more efficient in this setting, it should be interpreted with caution, as the multivariate ESS primarily reflects sampling efficiency and may fail to detect lack of convergence, as discussed in Section 4.3. [arias2025elliptical] report the efficiency in terms of time required to generate 1000 multivariate ESS. The values are 0.100 and 0.080 hours per 1000 ESS for the NUTS (including warmup) and the elliptical slice sampler, respectively. Following [arias2025elliptical], the multivariate ESS is calculated jointly for the first three columns of the impulse response parameters, corresponding to the identified shocks in this model.

### 21 May 2026 v2

@@ 4.2 Imposing Identifying Restrictions
In this section, we show how to impose identifying restrictions on the parameters using continuously differentiable transformations, which result in an unrestricted parameter vector. Our sampler runs on this vector, which means that rejection sampling is not needed to impose the identifying restrictions. Implementation of exclusion restrictions is straightforward in that no Jacobian adjustment is needed, whereas restrictions based on inequalities do require it as they are based on a change of variables. Because a different set of identifying restrictions is considered in each application, there are no general transformation and Jacobian that would cover all possible cases, but we discuss a few building blocks which can be used to construct frequently encountered restrictions.
@@
This is the density that our algorithm samples from. The exact cost of computing $|J|$ varies depending on the number and type of restrictions imposed on the model, but in applications involving sign restrictions or elasticity bounds it is negligible in comparison to the cost of evaluating the normal-inverse Wishart density. For example, if we restrict the sign of every parameter in $B$, evaluation of $\log|J|$ collapses to computing the sum over the $N^{2}$ elements of $\theta$ corresponding to $B$.
Let us consider the implementation of typical inequality restrictions. Suppose first that we have a density $p_{\alpha}(\alpha)$ over a parameter $\alpha\in\mathbb{R}_{+}$, which could, for example, be an impact response parameter from the matrix $B$. A Metropolis-type sampler running on $\alpha$ produces proposals $\alpha<0$, which are always rejected, and this decreases efficiency. Instead, we can reparameterize the sampler in terms of an unrestricted parameter, so that no a priori known regions of zero probability are encountered. To that end, we introduce a parameter $\tilde{\alpha}\in\mathbb{R}$, and set $\alpha=e^{\tilde{\alpha}}$. The Jacobian of this transformation is $e^{\tilde{\alpha}}$, and the corresponding density function is then $p(\tilde{\alpha})=p_{\alpha}(e^{\tilde{\alpha}})e^{\tilde{\alpha}}$ by a change of variables. If we want to restrict $\alpha<0$, we set $\alpha=-e^{\tilde{\alpha}}$, and the corresponding density is $p(\tilde{\alpha})=p_{\alpha}(-e^{\tilde{\alpha}})e^{\tilde{\alpha}}$.
Another useful transformation is the logit transformation for scalars with lower and upper bounds. Suppose we want to sample a parameter $\alpha\in(a,b)$, say, an elasticity parameter which needs to be positive ($a=0$) and also constrained from above. To that end, we define an $\tilde{\alpha}\in\mathbb{R}$, and obtain $\alpha$ as
@@
To illustrate these methods in practice, let us consider a two-variable model with a single lag and no constant. Suppose that the variables are quantity and price, and the supply and demand shocks are identified using sign restrictions on the impact responses of the variables. The responses have opposite signs for the supply shock and the same sign for the demand shock. If the first shock is the supply shock, the first column of the impact matrix $B$ thus contains one positive and one negative element, whereas the second column contains two positive elements.
@@ 4.3 Estimation Algorithm
Because the density function (9) does not facilitate direct sampling, we employ an MCMC method. Specifically, we use the No-U-Turn Sampler (NUTS) of [hoffman2014], which is a variant of the Hamiltonian Monte Carlo (HMC) algorithm. An HMC algorithm explores the target distribution by simulating Hamiltonian dynamics, which allows the sampler to construct long trajectories while yielding a high acceptance probability by utilizing the gradients of the target log-density. At a high level, the HMC sampler can be thought of as a Metropolis-Hastings algorithm with a complicated proposal distribution. Its performance depends on tunable parameters $\epsilon,L$ and $M$. Starting from the current draw, the HMC sampler simulates a trajectory by integrating Hamiltonian dynamics using $L$ steps of size $\epsilon$ with a mass matrix $M$, whose optimal choice is the inverse of the covariance matrix of the target posterior distribution [betancourt2018, Section 4.3]. The resulting endpoint is then used as a proposal.
The NUTS solves the issue of tuning $L$ by dynamically expanding the path taken within each iteration until it starts to make a U-turn towards the initial state. The implementation of the NUTS in Stan further provides routines for adapting the step size and metric matrix during warmup, and calculates the gradient of the target log-density by automatic differentiation. These properties reduce the scope of implementation error and ensure that sampling can be done efficiently in a wide variety of situations. Further details on the HMC sampler and NUTS can be found in Appendix C.
@@
To demonstrate that an HMC sampler can effectively explore the posterior under set identification, we revisit a simulation experiment that [kitagawa2025] considered to illustrate the opposite. This example does not involve a VAR model, but we adopt it because it features a known and simple posterior geometry, which allows for a transparent diagnosis of convergence issues. Specifically, we generate data from a normal distribution with mean $\mu=\sum_{i=1}^{k}\mu_{i}$ and variance unity, where the parameters $\mu_{i}$, $i=1,\dots,k$, are not identified for $k\geq 2$. We study the performance of Hamiltonian Monte Carlo in this setting for $k=2$ and k=1,000k=1{,}000, with the latter case reflecting the high-dimensional parameter spaces encountered in VAR applications. For $k=2$ and k=1,000k=1{,}000, we generate 1,000 observations from $N(2,1)$ and $N(10,1)$ distributions, respectively. The larger mean in the latter case reflects the greater number of additive components.
The left panel of Figure 1 presents a trace plot based on 12,000 draws for the case $k=2$, of which 2,000 are warmup draws. Following [kitagawa2025], the warmup draws are included. The right panel depicts the histograms of the obtained sample (without the warmup iterations) together with the analytical marginal densities, when the support of each $\mu_{i}$ is bounded between 10-10 and 10.444In the case $k=2$, the marginal density is proportional to $\Phi(\sqrt{n}(\mu_{i}+10-\bar{x}))-\Phi(\sqrt{n}(\mu_{i}-10-\bar{x}))$, where $\Phi$ is the standard normal CDF, $n$ is the sample size, and $\bar{x}$ is the sample mean. For n=1,000n=1{,}000 this density is approximately uniform. The plots indicate that the NUTS (as implemented in Stan) is able to traverse the entire posterior distribution of $\mu$ without apparent difficulties and produces draws that closely match the analytical marginal densities. This contrasts with the behavior documented by [kitagawa2025] for a standard HMC algorithm in the same set-identified setup, where the sampler failed to adequately explore the parameter space associated with $(\mu_{1},\mu_{2})$.
@@
The NUTS performs well even in a high-dimensional setting, as illustrated in Figure 2 that presents a scatter plot of the first two components of $\mu$ and the estimated autocorrelation functions for these two parameters when k=1,000k=1{,}000. The draws are dispersed without visible clustering, and the autocorrelations decay rapidly toward zero. Corresponding plots for other pairs of parameters look essentially identical. These results indicate that, even in this high-dimensional setting, the NUTS produces high-quality samples and effectively explores the full target distribution.
@@
As [bacchiocchi2025] point out, besides flat posterior regions, a related but distinct challenge can arise when the identified set is multimodal because separated regions of posterior mass are difficult for MCMC algorithms (including HMC) to explore effectively. However as noted by [kitagawa2025], such local identification arises only in the case of non-homogenous restrictions, that is, when a parameter is set to a non-zero value. As we preclude such restrictions, multimodality does not arise, and hence the sampler does not need to traverse between distinct modes. This absence of multimodality follows from the fact that the likelihood depends on $B$ only through $BB^{\prime}$ and is therefore invariant under orthogonal rotations, implying a continuum of observationally equivalent factorizations rather than isolated peaks. [^5] In this paper, we impose the normalization $\det(B)>0$ and restrict attention to a single normalized representation of the structural parameters to eliminate any remaining reflection symmetry. This normalization does not restrict the identified set beyond our existing identifying assumptions. To see this, consider the case where at most N1N-1 columns of $B$ are identified, so the last column is left unrestricted. Let $B_{1}$ satisfy the identifying restrictions and have $\det(B_{1})<0$. Define $R=\mathrm{diag}(1,\ldots,1,-1)$ and $B_{2}=B_{1}R$, i.e., $B_{2}$ equals $B_{1}$ with the sign of its last column flipped. Then $\det(B_{2})=\det(B_{1})\det(R)=-\det(B_{1})>0$, while the first N1N-1 columns of $B_{2}$ and $B_{1}$ are equal. Hence, $B_{2}$ satisfies the same identifying restrictions and has the same economic interpretation. Moreover, the map $B\mapsto BR$ is bijective with inverse $B\mapsto BR^{-1}=BR$, which guarantees that there always exists exactly one such $B_{2}$. As the preceding example indicates, the HMC sampler is well suited for exploring such distributions.
@@
The $\hat{R}$ measures chain mixing by comparing within- and between-chain variances. For a well-mixed chain, it approaches unity as the number of draws increases. When running a single chain (as is typical in the SVAR literature), $\hat{R}$ is calculated by splitting the chain into two halves. The idea of the bulk and tail ESS, in turn, is to measure sampling efficiency by adjusting the number of iterations for the autocorrelation in the sample. The bulk ESS estimates the number of effectively independent samples in the MCMC output for estimating the mean or the median of the posterior, while the tail ESS measures sampling efficiency in the tails of the distribution. [vehtari2021] recommend accepting an MCMC sample only if $\hat{R}<1.01$ and the ESS is at least 100 for a single chain. These thresholds ensure adequate mixing and a sufficiently low Monte Carlo standard error for estimation of the quantities of interest.
@@ 5 Empirical Applications
This section revisits two empirical applications to showcase the benefits of our approach: In Section 5.1, we identify three shocks in the small-scale global oil market model of [kilian2014], and in Section 5.2, we consider the large-scale U.S. economy model recently analyzed by [chan2025]. In each case, the empirical findings coincide with prior results based on the same identifying restrictions, but we highlight the advantages of our method over existing alternatives. Also [arias2025elliptical] consider these applications to illustrate the performance of their Gibbs sampler, which facilitates straightforward comparison. As their approach represents the current state of the art in inference under identifying sign restrictions, it provides a natural benchmark for our analysis. In both applications, our method achieves convergence approximately two to four times faster than theirs, as measured by the time required for the $\hat{R}$ statistic to fall below the conventional threshold. The obtained samples also exhibit lower autocorrelations, as reflected in greater bulk and tail ESS in both cases. In the large-scale model, our approach is also significantly faster than the algorithm of [chan2025], although they use the asymmetric conjugate prior [[, see]]chan2022 rather than the normal-inverse Wishart prior. In Section 5.2, we also show that our approach accommodates exclusion restrictions at no additional computational cost.
@@ 5.1 Small Oil Market Model
To illustrate how our approach handles contemporaneous and dynamic sign restrictions, as well as elasticity restrictions, we estimate the oil market model of [kilian2014] on monthly data from 1973:M2 to 2009:M8.777The data are obtained from the replication package for [kilian2014] in the Journal of Applied Econometrics Data Archive. It contains four variables, namely, the change in global oil production, a measure of global real activity, the real price of oil, and the changes in global oil inventories, as well as an intercept, and monthly dummy variables. Following [kilian2014], the lag length is set at 24. The prior is flat and proportional to $|\Sigma|^{-(N+1)/2}$, which results in a conjugate normal-inverse Wishart posterior for the reduced-form parameters. Following the identification strategy of [kilian2014], we identify the flow supply, flow demand, and speculative demand shocks. In particular, we impose the sign and elasticity restrictions on the impact responses, as well as the dynamic sign restrictions summarized in Table 1. These restrictions result in a tightly identified model, and thus random sampling of the orthogonal rotation matrix $Q$ would result in extremely few accepted posterior draws, as discussed in Section 3.
@@
where $\tilde{D}$ is a $4\times 11$ matrix of dummy coefficients, with each column corresponding to one dummy variable. These parameters are transformed into an unconstrained vector $\theta$ using the transformations described in section 4.2, specifically the exponential and inverse logit transformations on $B$ and $\Psi_{1},\dots,\Psi_{12}$. We run 2,000 warmup iterations, followed by 100,000 sampling iterations.
@@
We now turn to the properties of our sampler. Figure 4 presents trace plots of the unnormalized posterior log-density and a parameter ($B_{11}$) from the impact matrix $B$. For clarity, only the first 4,000 iterations, including warmup, are shown. As can be seen, both the log-density and the parameter are initiated near their typical level. This indicates that our method for selecting initial values by direct sampling works well and initializes the chain near the typical set. The traceplots do not exhibit problematic behavior such as nonstationarity or high autocorrelation after warmup. Furthermore, the variance of $B_{11}$ is lower during the early warmup phase than during sampling, which reflects the dynamic adaptation of the step size and inverse metric. This adaptation leads to more efficient sampling after warmup. The remaining structural parameters exhibit similar behavior and their traceplots are therefore not shown.
@@
To compare our approach with the elliptical slice sampler of [arias2025elliptical], we implemented their sampler in Matlab [^8] We thank Jonas Arias for generously sharing their replication package, which greatly facilitated our work. Building on this foundation, we implemented the elliptical slice sampler in R and Matlab, with the latter yielding a substantially faster implementation. We additionally experimented with implementing the large SVAR NUTS sampler of the next section in Matlab, which ran approximately one-third faster than the Stan version. Nonetheless, we base all reported results on Stan because Stan’s mature and widely tested implementation substantially reduces the scope for implementation errors and, through its automatic differentiation, obviates the need to derive and code gradients by hand, providing important practical and reliability advantages. and, following their lead, save every tenth draw from the resulting chain. We run this sampler for a total of four million iterations. To assess convergence, we use the $\hat{R}$ statistic (see Section 4.3), calculated from a single chain by considering samples of increasing length, with warmup iterations included. [^9] Because there is no clear way to ensure that the discarded warmup samples are comparable across samplers, we include the warmup iterations in the analysis. As recommended by [vehtari2021], we interpret values of $\hat{R}$ above 1.01 as an indication of potential convergence issues.
@@
Figure 5 plots the maximum values of the $\hat{R}$ statistic for samples obtained using both samplers. The reported sample size is increased in increments of 1,000 iterations. For the elliptical slice sampler, the maximum $\hat{R}$ is the maximum over the elements of $\Sigma$ and $Q$ , whereas for the NUTS it is the maximum over the elements of BB. [^10] In our experience, the unidentified parameters $B$ and $Q$ mix the most slowly. To reduce computational cost, we therefore do not compute $\hat{R}$ for all parameters. As can be seen in the figure, $\hat{R}$ converges in fewer iterations when using the NUTS. For the elliptical slice sampler, $\hat{R}$ first falls below the threshold after approximately 27 thousand iterations (corresponding to roughly 13 minutes), but requires about 130 thousand saved draws (64 minutes) to remain consistently below the threshold of 1.01. In contrast, the NUTS reaches the 1.01 threshold after roughly nine thousand iterations, corresponding to a runtime of 28 minutes.
@@
Figure 5: The maximum $\hat{R}$ statistic for samples of different sizes from the elliptical slice sampler and NUTS. Horizontal dashed lines are located at 1.00 and 1.01. $\hat{R}$ values are calculated for the elements of $B$ for the NUTS, and the elements of $\Sigma$ and $Q$ for the elliptical slice sampler. The vertical axis is restricted to $(1.0,1.1)$ for clarity.
@@
To investigate the sampling efficiency of the samplers, we consider the bulk and tail ESS measures of [vehtari2021]. In line with the $\hat{R}$, they are calculated for all elements of $B$ (for the NUTS), and $\Sigma$ and $Q$ (for the slice sampler), again using chains of increasing length and including warmup iterations. Table 2 presents the minimum ESS values per 1,000 iterations over these parameters for the NUTS and the elliptical slice sampler. As can be seen, the NUTS is more efficient in terms of the ESS per iteration under both measures. This is especially evident in the case of the bulk ESS, indicating that our approach yields more reliable estimates of posterior means for a given number of iterations. Importantly, given the lengths of the sampled chains, both methods reach ESS levels at which the Monte Carlo error is negligible for inference.
@@
Table 2: Minimum ESS measures per 1000 iterations. The values are calculated for elements of $B$ for the NUTS, and $\Sigma$ and $Q$ for the elliptical slice sampler.
ESS type
Sampler Bulk Tail
NUTS 40.53 48.24
Slice Sampler 2.18 5.66
@@
The non-monotone behavior of the $\hat{R}$ statistic and the lower ESS for the elliptical slice sampler are likely due to the high autocorrelation in the sampled chain. Figure 6 presents the autocorrelation function for $B_{11}$ from the NUTS (top panel), and $Q_{11}$ from the elliptical slice sampler (bottom panel). The latter exhibits substantially higher autocorrelation, which reduces sampling efficiency. Similar patterns are observed for the other elements of the matrices $B$ and $Q$ in this application.
We also assess convergence using the minimum multivariate ESS, as suggested by [vats2019]. To achieve a precision of 0.05 for the 95% confidence region our approach requires 52 minutes of sampling, whereas the elliptical slice sampler requires 31 minutes. Precision of 0.05 implies that the Monte Carlo error is 5% of the variability in target distribution. While this metric suggests that the elliptical slice sampler is more efficient in this setting, it should be interpreted with caution, as the multivariate ESS primarily reflects sampling efficiency and may fail to detect lack of convergence, as discussed in Section 4.3. [arias2025elliptical] report the efficiency in terms of time required to generate 1,000 multivariate ESS. The values are 4.76 and 5.17 minutes per 1,000 ESS for the NUTS (including warmup) and the elliptical slice sampler, respectively. Following [arias2025elliptical], the multivariate ESS is calculated jointly for the first three columns of the impulse response parameters, corresponding to the identified shocks in this model.